OpenStudy (anonymous):

The length of a diagonal square is 7 mm. Find the area of the square. A)24.5 B)29.08 C)21.2 D)22.1

Parth (parthkohli):

Since we have two 45-45-90 triangles, the length of one side will be \(7 \div \sqrt{2}\)

OpenStudy (anonymous):

@zooo : HELP ?

Parth (parthkohli):

You can rationalize it

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large {7 \over \sqrt{2}} = {7\sqrt{2} \over 2} }\) Now just square it :)

OpenStudy (anonymous):

WHATS THE SQUARE ROOT OF 2 ?

OpenStudy (anonymous):